OpenStudy (anonymous):

How to write this in Summation Notation : 3/2 - 5/4 +7/8 -9/16 +11/32 -13/64 ...

OpenStudy (anonymous):

Notice that the denominator (the easy part here) is going from: 2 to (2)(2) to (2)(2)(2)

OpenStudy (anonymous):

okk

OpenStudy (anonymous):

As for the numerator: 3 to -[3 + (2)(1)] to +[3 + (2)(2)] to -[3 + (2)(3)]

OpenStudy (anonymous):

So, you can use an index of starting at "1" and going on to infinity (because of the ellipses)

OpenStudy (anonymous):

You will have a factor of "alternating sign" in the numerator which will be represented by (-1) to an increasing power. With these hints, do you want to take a stab at it?

OpenStudy (anonymous):

\[\sum_{n=1}^{\infty}\frac{ (-1)^{n+1}[3 +(n-1)(2)] }{ 2^{n} }\]

OpenStudy (anonymous):

The first term is positive, so since your index starts at "1", the "n+1" will ensure a positive first term and alternating.

OpenStudy (anonymous):

ohh it will be \[\sum_{n=1}^{\infty} (1+2^n)/(2^n) (-1)^{n+1}\]

OpenStudy (anonymous):

The "n-1" in the second factor will give a "3" to start with. Just saw your work. Pretty close, but the (1 + 2^n) is off. Compare to what I posted for the summation.

OpenStudy (anonymous):

A good thing about these types of problems is that they are self-correcting. Just write out the first several terms from any answer and you'll see if it works or not. You might want to write out the first 4 or five terms from the summation I wrote out for you.

OpenStudy (anonymous):

Ah alright thanks!

OpenStudy (anonymous):

All good now? @butterflyprincess

OpenStudy (anonymous):

yes thnx!

OpenStudy (anonymous):

You're welcome! Good luck in all of your studies and thx for the recognition! @butterflyprincess